Consider the infinite series defined by "\u221e\ufeff\u2211n=1 ((2n!)\/(2^2n))".

(a) What is the value of r from the ratio test?

(b) What does this r value tell you about the series?


1
Expert's answer
2021-02-23T06:26:49-0500

Here we have the series "S_n" as,

"S_n= \displaystyle\sum_{n=1}^\infin\frac{(2n)!}{2^nn}"


(a)


Now, let "a_n=" "\frac{(2n)!}{2^nn}"


So, "a_{n+1}=\frac{(2n+2)!}{2^{n+1}(n+1)}"


So, "r=\frac{a_{n+1}}{a_n}"



"r=\frac{a_{n+1}}{a_n}\\~\\
\Rightarrow r=\frac{(2n+2)!\cdot n\cdot 2^n}{2^{n+1}\cdot(n+1)\cdot(2n)!}\\~\\
\Rightarrow r=n(n+1)"

Therefore, "r=n(n+1)"


(b)


Now, we will evaluate "\lim\limits_{n\to\infin}r"


So,


"\lim\limits_{n\to\infin}r\\
=\lim\limits_{n\to\infin}n(n+1)=\infin"

So, as "r" tends to "\infin" ,we can conclude that the series "S_n" diverges.